OpenStudy (anonymous):

solve the equation: x+sqrt(x-4)=4

OpenStudy (anonymous):

\[x+\sqrt{x-4}=4\]

Parth (parthkohli):

Subtract \(x\) from both sides. \( \color{Black}{\Rightarrow \sqrt{x - 4} = 4 - x}\) Square both sides.

OpenStudy (anonymous):

i dont get it wat do u mean can u give an example

Parth (parthkohli):

Where did you get confused?

OpenStudy (anonymous):

u said i should subtract x for both sides but how and how do i square both sides

Parth (parthkohli):

Hmm. You need Google to know more about isolation.